TIFFINFO(1) TIFFINFO(1) NAME tiffinfo−printinformationaboutTIFFfiles SYNOPSIS tiffinfo[options]input.tif... DESCRIPTION TiffinfodisplaysinformationaboutfilescreatedaccordingtotheTagImageFileFormat,Revision 6.0.Bydefault,thecontentsofeachTIFFdirectoryineachfileisdisplayed,withthevalueof eachtagshownsymbolically(wheresensible). OPTIONS −cDisplaythecolormapandcolor/grayresponsecurves,ifpresent. −DInadditiontodisplayingthedirectorytags,readanddecompressallthedataineach image(butnotdisplayit). −dInadditiontodisplayingthedirectorytags,printeachbyteofdecompresseddatainhexadecimal. −jDisplayanyJPEG-relatedtagsthatarepresent. −oSettheinitialTIFFdirectoryaccordingtothespecifiedfileoffset.Thefileoffsetmaybe specifiedusingtheusualC-stylesyntax;i.e.aleading‘‘0x’’forhexadecimalandaleading ‘‘0’’foroctal. −sDisplaytheoffsetsandbytecountsforeachdatastripinadirectory. −zEnablestripchoppingwhenreadingimagedata. −#SettheinitialTIFFdirectoryto#. SEEALSO pal2rgb(1),tiffcp(1),tiffcmp(1),tiffmedian(1),libtiff(3) January27,1997 1